OpenStudy (anonymous):

Solve the equations by substitution. If the equations are inconsistent and the solution is an empty set, write "no solution". 5 x - 2 y = 6 x = 5 - y x = y =

OpenStudy (anonymous):

Substitute x = 5 - y into the other equation instead of x: 5(5 - y) - 2 y = 6 and solve for y 25-5y-2y=6 19=7y y=19/7 Now use this value of y to find x by putting it in the 1º equation: x = 5 - 19/7=(35-19)/7=16/7